# Replace, Don’t Expand: Mitigating Context Dilution in Multi-Hop RAG via Fixed-Budget Evidence Assembly

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) systems often fail on multi-hop queries when the initial retrieval misses a bridge fact. Prior corrective approaches, such as Self-RAG, CRAG, and Adaptive-$k$, typically address this by adding more context or pruning existing lists. However, simply expanding the context window often leads to context dilution, where distractors crowd out relevant information. We propose SEAL-RAG, a training-free controller that adopts a “replace, don’t expand” strategy to fight context dilution under a fixed retrieval depth $k$. SEAL executes a (Search $\rightarrow$ Extract $\rightarrow$ Assess $\rightarrow$ Loop) cycle: it performs on-the-fly, entity-anchored extraction to build a live gap specification (missing entities/relations), triggers targeted micro-queries, and uses entity-first ranking to actively swap out distractors for gap-closing evidence. We evaluate SEAL-RAG against faithful re-implementations of Basic RAG, CRAG, Self-RAG, and Adaptive-$k$ in a shared environment on HotpotQA and 2WikiMultiHopQA. On HotpotQA ($k=3$), SEAL improves answer correctness by +3–13 pp and evidence precision by +12–18 pp over Self-RAG. On 2WikiMultiHopQA ($k=5$), it outperforms Adaptive-$k$ by +8.0 pp in accuracy and maintains 96% evidence precision compared to 22% for CRAG. These gains are statistically significant ($p<0.001$). By enforcing fixed-$k$ replacement, SEAL yields a predictable cost profile while ensuring the top-$k$ slots are optimized for precision rather than mere breadth. We release our code and data at https://github.com/mosherino/SEAL-RAG.
@@ 1 Introduction
To mitigate this, recent research has focused on iterative and corrective RAG. Systems like Self-RAG [asai2024selfrag] and CRAG [yan2024crag] introduce feedback loops: they critique the retrieved evidence and trigger additional retrieval steps if gaps are detected. While effective at improving recall, these methods typically operate via Breadth-First Addition: they append new passages to the existing context. This approach assumes that “more context is better,” but in production environments with strict latency and token budgets, this assumption fails. Expanding the context window introduces distractors (irrelevant passages that confuse the model) and lateral redundancy (duplicate information), often degrading the model’s ability to reason over the specific bridge facts required—a phenomenon known as context dilution [liu2023lostinthemiddle].
@@ 1.3 Paper Organization
Section˜2 reviews related work. Section˜3 details SEAL-RAG (loop controller, scope-aware sufficiency, loop-adaptive extraction, entity-first ranking, and the micro-query policy). Section˜4 specifies datasets, models, retrieval/indexing, baseline, metrics/judging, and protocol. Section˜5 presents main results at $k=1$, $k=3$, and $k=5$ with per-backbone tables and discussion. Section˜6 reports loop-budget ablations and analysis. Section˜7 states limitations and threats to validity. Section˜7 concludes. Detailed prompts and statistical tables appear in the Appendix. All code and datasets are available in the GitHub repository at https://github.com/mosherino/SEAL-RAG.
@@ 2.1 Standard RAG and Multi-Hop Challenges
Retrieval-Augmented Generation (RAG) has evolved from early sparse retrieval pipelines to sophisticated dense and hybrid systems [lewis2020rag, gao2023survey]. While dense retrievers like DPR [karpukhin2020dpr] and late-interaction models like ColBERT [khattab2020colbert] improve recall on single-hop queries, they often struggle with multi-hop reasoning, where the answer depends on composing information from multiple disjoint documents [min2019compositional, chen2019understanding]. In these scenarios, the standard retrieve-then-generate pattern faces a critical bottleneck: if the initial top-$k$ set misses a “bridge” fact, the generator cannot recover. A common workaround is to blindly increase $k$ or accumulate more context, but this introduces noise. Recent analysis confirms that irrelevant context can significantly degrade model performance—a phenomenon known as “lost in the middle” or context dilution [liu2023lostinthemiddle]. SEAL-RAG targets this specific failure mode by holding $k$ fixed and iteratively repairing the evidence set rather than expanding it.
@@ 2.3 Adaptive Retrieval and Pruning
A parallel line of work focuses on dynamic resource allocation to improve efficiency. Adaptive-RAG [jeong2024adaptive] functions as a router, classifying query complexity to dynamically select between retrieval-free and retrieval-augmented paths. Adaptive-$k$ [taguchi2025adaptivek] and LC-Boost [wu2024lcboost] aim to optimize the context window by pruning irrelevant documents from a larger retrieved list or selecting a minimal sufficient subset. While these methods address the efficiency drawback of standard RAG, they are primarily selectors or routers, not repairers. If the initial retrieval pool misses a bridge fact entirely, pruning cannot recover it. In contrast, SEAL-RAG performs Active Repair: it diagnoses specific missing entities (e.g., via on-the-fly extraction) and issues targeted micro-queries to fetch new evidence that was never in the initial pool, replacing low-utility items to improve the set’s composition.
@@ 4.3 Baselines
  • CRAG (Corrective RAG): A corrective graph [yan2024crag]. If retrieved documents are graded as “irrelevant,” the system triggers an external web search (via Tavily) to augment the context before generation.
@@ 5.2 Main Results on HotpotQA ($k=3$)
Table 2: Main results at fixed $k=3$ on HotpotQA (N=1, [^0] N=1{,}000). Even with larger capacity, SEAL-RAG maintains a significant lead in Precision and Correctness. $\Delta$ shows the gain over the best baseline.
@@ 5.3 Generalization to 2WikiMultiHopQA ($k=1,3,5$)
Table 3: 2WikiMultiHopQA Results ($N=200$). We compare performance across retrieval depths. Key Trend: As $k$ increases to 5, baseline Precision collapses (Context Dilution), while SEAL-RAG maintains high precision via replacement, driving superior Accuracy (Judge-EM).
@@
Table 4: SEAL-RAG vs. Adaptive-$k$ ($N=200$). Adaptive-$k$ acts as a selector: it is bound by the quality of the initial pool. SEAL-RAG acts as a repairer: it actively fetches missing information, breaking the ceiling of the initial retrieval.
@@ 5.5 Statistical Significance Analysis
  • Methodology: For binary answer correctness (Judge-EM), we used McNemar’s test, which is appropriate for paired nominal data [dror2018hitchhiker]. For continuous retrieval metrics (Precision/Recall/F1), we used paired two-sided $t$-tests [koehn2004statistical]. We applied the Holm-Bonferroni correction to control the family-wise error rate at $\alpha=0.05$.
@@ E.1 Methodology & Alignment
  • Binary Metrics (Judge-EM): We used McNemar’s test with the chi-squared approximation ($N=1000$). The statistic $\chi^{2}$ compares the discordant pairs.
@@
Table 7: Significance vs. Adaptive-$k$ (2Wiki, $k=5$). SEAL-RAG significantly outperforms both variants in Accuracy and Precision.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems often fail on multi-hop queries when the initial retrieval misses a bridge fact. Prior corrective approaches, such as Self-RAG, CRAG, and Adaptive-$k$, typically address this by adding more context or pruning existing lists. However, simply expanding the context window often leads to context dilution, where distractors crowd out relevant information. We propose SEAL-RAG, a training-free controller that adopts a “replace, don’t expand” strategy to fight context dilution under a fixed retrieval depth $k$. SEAL executes a (Search $\rightarrow$ Extract $\rightarrow$ Assess $\rightarrow$ Loop) cycle: it performs on-the-fly, entity-anchored extraction to build a live gap specification (missing entities/relations), triggers targeted micro-queries, and uses entity-first ranking to actively swap out distractors for gap-closing evidence. We evaluate SEAL-RAG against faithful re-implementations of Basic RAG, CRAG, Self-RAG, and Adaptive-$k$ in a shared environment on HotpotQA and 2WikiMultiHopQA. On HotpotQA ($k=3$), SEAL improves answer correctness by +3 to +13 pp and evidence precision by +12 to +18 pp over Self-RAG. On 2WikiMultiHopQA ($k=5$), it outperforms Adaptive-$k$ by +8.0 pp in accuracy and maintains 96% evidence precision, compared to 22% for CRAG. These gains are statistically significant ($p<0.001$). By enforcing fixed-$k$ replacement, SEAL yields a predictable cost profile while ensuring the top-$k$ slots are optimized for precision rather than mere breadth. We release our code and data at https://github.com/mosherino/SEAL-RAG.
Keywords: Retrieval-Augmented Generation; multi-hop reasoning; context dilution; evidence assembly; SEAL-RAG; HotpotQA; 2WikiMultiHopQA
@@ 1 Introduction
To mitigate this, recent research has focused on iterative and corrective RAG. Systems like Self-Reflective RAG (Self-RAG) [asai2024selfrag] and Corrective RAG (CRAG) [yan2024crag] introduce feedback loops: they critique the retrieved evidence and trigger additional retrieval steps if gaps are detected. While effective at improving recall, these methods typically operate via Breadth-First Addition: they append new passages to the existing context. This approach assumes that “more context is better,” but in production environments with strict latency and token budgets, this assumption fails. Expanding the context window introduces distractors (irrelevant passages that confuse the model) and lateral redundancy (duplicate information), often degrading the model’s ability to reason over the specific bridge facts required—a phenomenon known as context dilution [liu2023lostinthemiddle].
@@ 1.3 Paper Organization
Section˜2 reviews related work. Section˜3 details SEAL-RAG (loop controller, scope-aware sufficiency, loop-adaptive extraction, entity-first ranking, and the micro-query policy). Section˜4 specifies datasets, models, retrieval/indexing, baseline, metrics/judging, and protocol. Section˜5 presents main results at $k=1$, $k=3$, and $k=5$ with per-backbone tables and discussion. Section˜6 reports loop-budget ablations and analysis. Section˜7 states limitations and threats to validity. Section˜8 concludes the paper. Detailed prompts and statistical tables appear in the Appendix. All code and datasets are available in the GitHub repository at https://github.com/mosherino/SEAL-RAG.
@@ 2.1 Standard RAG and Multi-Hop Challenges
RAG has evolved from early sparse retrieval pipelines to sophisticated dense and hybrid systems [lewis2020rag, gao2023survey]. While dense retrievers like Dense Passage Retrieval (DPR) [karpukhin2020dpr] and late-interaction models like ColBERT [khattab2020colbert] improve recall on single-hop queries, they often struggle with multi-hop reasoning, where the answer depends on composing information from multiple disjoint documents [min2019compositional, chen2019understanding]. In these scenarios, the standard retrieve-then-generate pattern faces a critical bottleneck: if the initial top-$k$ set misses a “bridge” fact, the generator cannot recover. A common workaround is to blindly increase $k$ or accumulate more context, but this introduces noise. Recent analysis confirms that irrelevant context can significantly degrade model performance—a phenomenon known as “lost in the middle” or context dilution [liu2023lostinthemiddle]. SEAL-RAG targets this specific failure mode by holding $k$ fixed and iteratively repairing the evidence set rather than expanding it.
@@ 2.3 Adaptive Retrieval and Pruning
A parallel line of work focuses on dynamic resource allocation to improve efficiency. Adaptive Retrieval-Augmented Generation (Adaptive-RAG) [jeong2024adaptive] functions as a router, classifying query complexity to dynamically select between retrieval-free and retrieval-augmented paths. Adaptive-$k$ [taguchi2025adaptivek] and Long-Context Bootstrapper (LC-Boost) [wu2024lcboost] aim to optimize the context window by pruning irrelevant documents from a larger retrieved list or selecting a minimal sufficient subset. While these methods address the efficiency drawback of standard RAG, they are primarily selectors or routers, not repairers. If the initial retrieval pool misses a bridge fact entirely, pruning cannot recover it. In contrast, SEAL-RAG performs Active Repair: it diagnoses specific missing entities (e.g., via on-the-fly extraction) and issues targeted micro-queries to fetch new evidence that was never in the initial pool, replacing low-utility items to improve the set’s composition.
@@ 4.3 Baselines
  • CRAG: A corrective graph [yan2024crag]. If retrieved documents are graded as “irrelevant,” the system triggers an external web search (via Tavily) to augment the context before generation.
@@ 5.2 Main Results on HotpotQA ($k=3$)
Table 2: Main results at fixed $k=3$ on HotpotQA (N=1, [^0] N=1{,}000). Even with larger capacity, SEAL-RAG maintains a significant lead in Precision and Correctness. $\Delta$ shows the gain over the best baseline ($p<0.001$).
@@ 5.3 Generalization to 2WikiMultiHopQA ($k=1,3,5$)
Table 3: 2WikiMultiHopQA Results ($N=200$). We compare performance across retrieval depths ($k$). Key Trend: As $k$ increases to 5, baseline Precision collapses (Context Dilution), while SEAL-RAG maintains high precision via replacement, driving superior Accuracy (Judge-EM).
@@
Table 4: SEAL-RAG vs. Adaptive-$k$ ($N=200$) on the 2WikiMultiHopQA dataset. Adaptive-$k$ acts as a selector: it is bound by the quality of the initial pool. SEAL-RAG acts as a repairer: it actively fetches missing information, breaking the ceiling of the initial retrieval.
@@ 5.5 Statistical Significance Analysis
  • Methodology: For Judge-EM, we used McNemar’s test, which is appropriate for paired nominal data [dror2018hitchhiker]. For continuous retrieval metrics (Precision/Recall/F1), we used paired two-sided $t$-tests [koehn2004statistical]. We applied the Holm-Bonferroni correction to control the family-wise error rate at $\alpha=0.05$.
@@ 8 Conclusion
Future work will focus on optimizing the latency of the iterative controller, potentially via distilling the sufficiency gate into smaller, specialized models to reduce overhead. Additionally, we aim to explore the applicability of fixed-budget assembly in high-stakes domain-specific expert systems, such as legal or medical retrieval, where evidence precision is paramount.
@@ E.1 Methodology & Alignment
  • Judge-EM: We used McNemar’s test with the chi-squared approximation ($N=1000$). The statistic $\chi^{2}$ compares the discordant pairs.
@@
Table 7: Significance vs. Adaptive-$k$ (2Wiki, $k=5$). SEAL-RAG significantly outperforms both variants in Accuracy and Precision ($p<0.001$).